OpenStudy (anonymous):

√(5) × √(10) simplifying radicals

OpenStudy (whpalmer4):

Remember that \[\sqrt{a}*\sqrt{b} = \sqrt{a*b}\]Then you can factor the value of a*b and see if there are any duplicated factors (perfect squares, in other words). For example: \[\sqrt{2}\sqrt{8} = \sqrt{16} = \sqrt{(2*2)*(2*2)} = 2*2 = 4\]